THE SPECTRUM OF SINGLET SILYLENE

\underline{S. N. YURCHENKO} and P. R. BUNKER, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6; W. P. KRAEMER, Max Planck Institut für Astrophysik, Karl-Schwarzschild-Strasse 1, Postfach 1523, D-85740 Garching, Germany; PER JENSEN, Theoretische Chemie, Fachbereich 9, Bergische Universität, D–42097 Wuppertal, Germany.

We report the results of ab initio and rovibronic energy calculations for the two electronic states of silylene (SiH_{2}), X̃^{1}A_{1} and Ã^{1}B_{1}, that become degenerate when the molecule is linear. Potential energy surfaces, and transition dipole moment functions, are calculated using high level ab initio methods. The potential function parameters are optimized by fitting to experimental energies, but in the fitting the shapes of the potential surfaces are constrained by the ab initio results. The variational RENNER program is used to calculate rotation-vibrational energies, and to simulate the X̃^{1}A_{1}–Ã^{1}B_{1} emission spectrum of SiH_{2}. Local mode assignments within the X̃–state polyad blocks are made, and these are important in the fitting process; they are also important for the predictions.